Sally's age plus 8 is equal to Garts age doubled. The sum of their ages is 124. How old is Sally?
Answer by macston(5194)

G=garts age; S=Sally's age=2G-8
S+G=124 Substitute for S
2G-8+G=124
3G=132
G=44
S=2G-8=2(44)-8=88-8=80
ANSWER: Sally's age is 80.